OpenStudy (anonymous):

The area of a circle is 81 pie cm2. What is the radius of this circle??

OpenStudy (anonymous):

\[\huge \color{green}{Area = \pi r^2}\]

OpenStudy (anonymous):

To find radius use this one: \[\huge \color{blue}{r = \sqrt{\frac{Area}{\pi}}}\] \[\huge \color{blue} {\color{red} {\checkmark} Area = 81 \pi}\]

OpenStudy (anonymous):

So the answer would be 9. Because 9 squared is 81! 9 x 9=81

OpenStudy (anonymous):

Yes dear you are right..

OpenStudy (anonymous):